# CoSy: Evaluating Textual Explanations of Neurons

### v1

@@ Abstract
A crucial aspect of understanding the complex nature of Deep Neural Networks (DNNs) is the ability to explain learned concepts within their latent representations. While various methods exist to connect neurons to textual descriptions of human-understandable concepts, evaluating the quality of these explanation methods presents a major challenge in the field due to a lack of unified, general-purpose quantitative evaluation. In this work, we introduce CoSy (Concept Synthesis)—a novel, architecture-agnostic framework to evaluate the quality of textual explanations for latent neurons. Given textual explanations, our proposed framework leverages a generative model conditioned on textual input to create data points representing the textual explanation. Then, the neuron’s response to these explanation data points is compared with the response to control data points, providing a quality estimate of the given explanation. We ensure the reliability of our proposed framework in a series of meta-evaluation experiments and demonstrate practical value through insights from benchmarking various concept-based textual explanation methods for Computer Vision tasks, showing that tested explanation methods significantly differ in quality. We provide an open-source implementation on GitHub [^1] https://github.com/lkopf/cosy.
Introduction
@@
One of the key obstacles to the wider adoption of Machine Learning methods in various areas is the inherent opacity of modern Deep Neural Networks (DNNs)—in simple terms, we do not understand why these machines make the predictions they do. To address this problem, the field of Explainable AI (XAI) [samek2019towards, xu2019explainable] has emerged, to reveal the decision-making processes of DNNs in a human-understandable fashion. XAI has broadened its focus from explaining the decision-making of DNNs locally, i.e., for specific inputs using saliency maps [bach2015pixel, simonyan2014deep, selvaraju2017grad, smilkov2017smoothgrad], to explaining the global behavior of the models by analyzing individual model components and their functional purpose [olah2020zoom]. Following the latter global explainability approach, often referred to as mechanistic interpretability [wang2022interpretability, bills2023language, nanda2022progress], there are methods that aim to describe the specific concepts neurons have learned to detect [Bau_2017_CVPR, muCompositionalExplanationsNeurons2021, hernandez2022natural, pmlr-v202-kalibhat23a, oikarinen2023clip, bykov2023labeling], enabling analysis of how these high-level concepts influence network predictions.
@@
A popular approach for explaining the functionality of latent representations of a network is to label neurons using human-understandable textual concepts. A textual description is assigned to a neuron based on the concepts that the neuron has learned to detect or is significantly activated by. Over time, these methods have evolved from providing label-specific descriptions [Bau_2017_CVPR] to more complex compositional [muCompositionalExplanationsNeurons2021, bykov2023labeling] and open-vocabulary explanations [hernandez2022natural, oikarinen2023clip]. However, a significant challenge remains: the lack of a universally accepted quantitative evaluation measure for open-vocabulary neuron descriptions. As a consequence, different methods devised their own evaluation criteria, making it difficult to perform general-purpose, comprehensive cross-comparisons.
@@
Figure 1: A schematic illustration of the CoSy evaluation framework for Neuron 80 in ResNet18’s avgpool layer. The current challenge lies in the absence of dataset- and architecture-agnostic evaluation measures to benchmark textual explanations of neurons. To address this, we propose CoSy, a framework consisting of three steps: first, a generative model translates textual concepts into the visual domain, creating synthetic images for each explanation using a text-to-image model. Then, inference is performed on synthetic images, along with a control image dataset, to collect neuron activations. Finally, by comparing the activations on synthetic images with activations on the control dataset, we can quantitatively assess the quality of the textual explanation and compare the results between different explanation methods. The implementation details of this example can be found in Appendix A.2.
@@
With our work, we aim to bridge this gap by introducing a novel quantitative evaluation framework named CoSy, for evaluating open-vocabulary explanations for neurons in Computer Vision (CV) models (illustrated in Figure 1). Our approach builds on recent advancements in Generative AI, which enable the generation of synthetic images that align with provided concept-based textual explanations. We use a set of available text-to-image models to synthesize data points that are prototypical for specific target explanations. These data points allow us to evaluate how neurons differentiate between concept-related images and those from a control dataset. We summarize our contributions as below:
@@
    In a series of meta-evaluation experiments (Section 4), we analyze the choice of generative models and prompts for synthetic image generation, demonstrating framework reliability.
@@
Activation Maximization is a commonly used methodology to understand what a neuron has learned to detect [erhan2009visualizing]. Such methods work by identifying input signals that trigger the highest activation in a neuron. This can be achieved synthetically, where an optimization process is employed to create the optimal input that maximizes the neuron’s activation [olah2017feature, nguyen2016synthesizing, fel2023unlocking], or naturally, by finding such inputs within a data corpus [borowski2020natural]. Activation Maximization has been employed for explaining latent representations of models [goh2021multimodal, yoshimura2021toward], including probabilistic models [grinwald2023visualizing], detection of backdoor attacks [casper2023red] and spurious correlations [bykov2023dora]. Recently, it was shown that Activation Maximization methods could be manipulated to illustrate predetermined signals without significantly affecting performance [bareeva2024manipulating, geirhos2023don, marty2022adversarial]. However, one of the key limitations of this methodology lies in its inability to scale; its scalability is limited due to its dependency on users to manually audit maximization signals.
@@
A more scalable alternative approach involves linking neurons with human-understandable concepts through textual descriptions. Network Dissection [Bau_2017_CVPR] (NetDissect) is a pioneering method in this field, associating convolutional neurons with a concept based on the Intersection over Union (IoU) of neuron activation maps and ground truth segmentation masks. Building on this, Compositional Explanations of Neurons (CompExp) [muCompositionalExplanationsNeurons2021] enhanced the detail of the explanations by allowing compositional concepts (i.e., concepts constructed using logical operators). MILAN [hernandez2022natural] further expanded this by allowing for open-vocabulary explanations, permitting the generation of descriptions beyond predefined labels. INVERT [bykov2023labeling] adopted a compositional concept approach, enabling explanations for general neuron types without relying on segmentation masks, and assigns compositional labels based on a neuron’s ability to distinguish concepts using the Area Under the Receiver Operating Characteristic Curve (AUC). FALCON [pmlr-v202-kalibhat23a] and CLIP-Dissect [oikarinen2023clip] compute image-text similarity with a CLIP model [radford2021learning] for the most activating images and their corresponding captions or concept sets. An overview of the different techniques is illustrated in Table 1. More detailed descriptions of these methods can be found in Appendix A.1.
@@
While significant effort has been made towards developing approaches and tools for evaluating local explanations [agarwal2022openxai, hedstrom2022quantus, hedstroem2023sanity], there has been relatively limited focus on evaluating global methods. Currently, to the best of our knowledge, there is no unified approach that allows for benchmarking across models and explanation methods. In their respective papers, the INVERT and CLIP-Dissect explanation methods evaluated the accuracy of their explanations by comparing the generated neuron labels with ground truth descriptions provided for neurons in the output layer of a network. CLIP-Dissect additionally evaluates the quality of explanations by computing the cosine similarity in a sentence embedding space between the ground truth class name for each neuron and the explanation generated by the method. FALCON employs a human study conducted on Amazon Mechanical Turk to evaluate the concepts generated by the method. Participants are tasked with selecting the best explanation for each target feature from a selection of explanation methods, considering a given set of highly and lowly activating images. MILAN evaluates the performance of neuron labeling methods relative to human annotations using BERTScores [zhang2019bertscore].
Method
@@
Table 1: Comparison of characteristics of concept-based textual explanation methods. The columns (from left to right) represent the explanation method used, its textual output type (fixed-label, compositional, or open-vocabulary), the type of neuron targeted for analysis (convolutional, scalar, or predetermined), the target metric the method optimizes (IoU, WPMI, AUC, etc.), whether the method relies on auxiliary black-box models for finding or generating explanations (img2txt, CLIP), and whether the explanation method is architecture-agnostic, meaning it can be applied to any CV model. For a more detailed description of each method, refer to Appendix A.1.
@@
NetDissect [Bau_2017_CVPR] fixed-label conv. IoU
CompExp [muCompositionalExplanationsNeurons2021] compositional conv. IoU
MILAN [hernandez2022natural] open-vocabulary conv. WPMI img2txt model
INVERT [bykov2023labeling] compositional scalar AUC
CLIP-Dissect [oikarinen2023clip] open-vocabulary scalar SoftWPMI CLIP
FALCON [pmlr-v202-kalibhat23a] open-vocabulary predetermined avg. CLIP score CLIP
@@ 3.1 Preliminaries
Consider a Deep Neural Network (DNN) represented by the function $g:\mathcal{X}\rightarrow\mathcal{Z},$ where $\mathcal{X}\subset\mathbb{R}^{h\times w\times c}$ denotes the input image domain and $\mathcal{Z}\subset\mathbb{R}^{l}$ represents the model’s output domain. We can view the model as a composition of two functions, $F:\mathcal{X}\rightarrow\mathcal{Y},$ and $L:\mathcal{Y}\rightarrow\mathcal{Z},$ such that $g=L\circ F.$ Here $\mathcal{Y}\subset\mathbb{R}^{k\times w^{*}\times h^{*}}$, where $k\in\mathbb{N}$ is the number of neurons in the layer, and $w^{*},h^{*}\in\mathbb{N}$ represent the width and height of the feature map, respectively. The function $F,$ which we refer to as the feature extractor, can be chosen based on the layer of the model we aim to inspect. This could be an existing layer within the model or a concept bottleneck layer [yuksekgonul2022post]. We refer to the $i$-th neuron within the layer as $f_{i}(\bm{x})=F_{i}(\bm{x}):\mathcal{X}\rightarrow\mathbb{R}^{w^{*}\times h^{*}}.$ Within the scope of this paper, we refer to explanation method as an operator $\mathcal{E}$ that maps a neuron to the textual description $s=\mathcal{E}(f_{i})\in\mathcal{S},$ where $\mathcal{S}$ is a set of potential textual explanations. The specific set of explanations depends on the implementation of the particular method.
@@ 3.2 CoSy: Evaluating Open-Vocabulary Explanations
A good textual explanation for a neuron should be a human-understandable description of an input that yields a high level of activation in the neuron. However, modern methods for explaining the functional purpose of neurons often provide open-vocabulary textual explanations, complicating the quantitative collection of natural data that represents the explanation. To address this issue, CoSy leverages recent advancements in generative models to synthesize data points that correspond to the textual explanation. The response of a neuron to a set of synthetic images is measured and compared to the neuron’s activation on a set of control natural images representing random concepts. This comparison allows for a quantitative evaluation of the alignment between the explanation and the target neuron.
Parameters of the proposed method include a control dataset $\mathbb{X}_{0}=\{\bm{x}^{0}_{1},\dots,\bm{x}^{0}_{n}\}\subset\mathcal{X},n\in\mathbb{N}$ – containing natural images that represent the concepts the model was originally trained on, a generative model $p_{M},$ that is used for synthesizing images, and a number of generated images $m\in\mathbb{N}.$ Given a neuron $f$ and explanation $s\in\mathcal{S},$ CoSy evaluates the alignment between the explanation and a neuron in 3 consecutive steps, which are illustrated in Figure 1.
@@
    $\mathbb{A}_{0}=\{\sigma(f(\bm{x}^{0}_{1})),\dots,\sigma(f(\bm{x}^{0}_{n}))\}\in\mathbb{R}^{n},\mathbb{A}_{1}=\{\sigma(f(\bm{x}^{1}_{1})),\dots,\sigma(f(\bm{x}^{1}_{m}))\}\in\mathbb{R}^{m},$ (1)
    where $\sigma$ is an aggregation function for multi-dimensional neurons. Within the scope of our paper, we use Average Pooling as aggregation function
    $\sigma(\bm{y})=\frac{1}{w^{*}h^{*}}\sum_{i\in[1,w^{*}]}\sum_{i\in[1,h^{*}]}\bm{y}_{i,j},\quad\bm{y}\in\mathcal{Y}\subset\mathbb{R}^{w^{*}\times h^{*}}.$ (2)
@@
    $\Psi_{\text{MAD}}(\mathbb{A}_{0},\mathbb{A}_{1})=\frac{1}{m}\sum_{b\in\mathbb{A}_{1}}b-\frac{1}{n}\sum_{a\in\mathbb{A}_{0}}a.$ (4)
@@
Meta-Evaluation Analysis
Meta-evaluation is the practice of evaluating the evaluation method itself [hedstrom2023metaquantus]. This process is crucial to ensure the reliability of our proposed evaluation measure. In this section, we analyze the following: (1) which generative models and prompts provide the best similarity to natural images, (2) whether the model’s behavior on synthetic and natural images differs for the same concept, and (3) validating that CoSy provides appropriate evaluation scores for true and random explanations, given known ground truth concept for the neuron.
@@ 4.1 Synthetic Image Reliability
One of the key features of CoSy is its reliance on generative models to translate textual explanations of neurons into the visual domain. Thus, it is essential that the generated images reliably resemble the textual concepts. In the following section, we present an experiment where we varied several parameters of the generation procedure and evaluated the visual similarity between generated images and synthetic ones, focusing on concepts for which we have a collection of natural images.
For our analysis, we used only open-source and freely available text-to-image models, namely Stable Diffusion XL 1.0-base (SDXL) [podell2023sdxl] and Stable Cascade (SC) [pernias2023wurstchen]. We also varied the prompts for image generation. To measure the similarity between synthetic images and natural images corresponding to the same concept, we employed cosine similarity (CS) in the CLIP embedding space with the CLIP-ViT-B/32 model [radford2021learning]. We select a set of 10 random concepts from the 1,000 classes in the ImageNet validation dataset [russakovsky2015imagenet]. For each [concept] we use 5 different prompts and employ them with SDXL and SC models, generating 50 images per concept. We then measure the CS between image pairs of the same class.
Figure 2: An overview of the impact of varying the prompt on the similarity between natural and synthetic images, using two text-to-image models. Left: average Cosine similarity (CS) across all natural and synthetic images over all classes are reported. Higher CS values are better, indicating greater similarity between the images. Right: an illustration of the visual differences produced by the SDXL and SC models in response to diverse prompts for the explanation concept “submarine”, and natural images from the ImageNet validation set. Our results show that SDXL and SC generate similar images, with SDXL generally being more closely aligned with natural images than SC.
Figure 2 illustrates the comparison across all generative models and prompts in terms of CS of generated images to natural images of the same class. The results indicate that when using prompt 5 as input to SDXL, the synthetic images show the highest similarity to natural images. The performance is generally best with the most detailed prompt (5) and closely aligns with prompts 1, 3, and 4. Moreover, SDXL appears to be slightly more effectively realizing detailed prompts than SC. As anticipated, the poorly constructed prompt (2) results in the lowest similarity to natural images for both models. If not stated otherwise, for all following experiments, prompt 5 together with SDXL model was employed for image generation.

### v2

@@ Abstract
A crucial aspect of understanding the complex nature of Deep Neural Networks (DNNs) is the ability to explain learned concepts within their latent representations. While methods exist to connect neurons to human-understandable textual descriptions, evaluating the quality of these explanations is challenging due to the lack of a unified quantitative approach. We introduce CoSy (Concept Synthesis), a novel, architecture-agnostic framework for evaluating textual explanations of latent neurons. Given textual explanations, our proposed framework uses a generative model conditioned on textual input to create data points representing the explanations. By comparing the neuron’s response to these generated data points and control data points, we can estimate the quality of the explanation. We validate our framework through sanity checks and benchmark various neuron description methods for Computer Vision tasks, revealing significant differences in quality. We provide an open-source implementation on GitHub [^1] https://github.com/lkopf/cosy.
1 Introduction
@@
One of the key obstacles to the wider adoption of Machine Learning methods across various fields is the inherent opacity of modern Deep Neural Networks (DNNs)—in essence, we often lack an understanding of why these models make certain predictions. To address this problem, various explainability methods [1, 2] have been developed to make the decision-making processes of DNNs more understandable to humans. Explainability methods have broadened their focus from interpreting the decision-making of DNNs locally—for instance, interpreting specific inputs using saliency maps [3, 4, 5, 6]—to understanding the global behavior of models by analyzing individual model components and their functional purpose [7]. Following the latter global explainability approach, often referred to as mechanistic interpretability [8, 9, 10], some methods aim to describe the specific concepts neurons have learned to detect [11, 12, 13, 14, 15, 16], enabling analysis of how these high-level concepts influence network predictions.
@@
A popular approach for explaining the functionality of latent representations of a network is to describe neurons using human-understandable textual concepts. A textual description is assigned to a neuron based on the concepts that the neuron has learned to detect or is significantly activated by. Over time, these methods have evolved from providing label-specific descriptions [11] to more complex compositional [12, 16] and open-vocabulary explanations [13, 15]. However, a significant challenge remains: the lack of a universally accepted quantitative evaluation measure for open-vocabulary neuron descriptions. As a consequence, different methods devised their own evaluation criteria, making it difficult to perform general-purpose, comprehensive cross-comparisons.
@@
Figure 1: A schematic illustration of the CoSy evaluation framework for Neuron 80 in ResNet18’s avgpool layer. The current challenge lies in the absence of general-purpose, quantitative evaluation measures to benchmark textual explanations of neurons. To address this, we propose CoSy, a framework consisting of three steps: first, a generative model translates textual concepts into the visual domain, creating synthetic images for each explanation using a text-to-image model. Then, inference is performed on these synthetic images alongside a control image dataset to collect neuron activations. Finally, by comparing activations from the synthetic images with those from the control dataset, we quantitatively assess the quality of the textual explanation and compare results across different explanation methods. The implementation details of this example can be found in Appendix A.2.
@@
In this work, we aim to bridge this gap by introducing Concept Synthesis (CoSy), the first automatic evaluation framework for textual explanations of neurons in Computer Vision (CV) models (illustrated in Figure 1). Our approach builds on recent advancements in Generative AI, which enable the generation of synthetic images that align with provided neuron explanations. We use a set of available text-to-image models to synthesize data points that are prototypical for specific target explanations. These data points allow us to evaluate how neurons differentiate between concept-related images and non-concept-related images combined in a control dataset. We summarize our contributions as below:
@@
    In a series of sanity checks (Section 4), we analyze the choice of generative models and prompts for synthetic image generation, demonstrating framework reliability.
@@ 2 Related Works
Activation Maximization is a commonly used methodology to understand what a neuron has learned to detect [17]. Such methods work by identifying input signals that trigger the highest activation in a neuron. This can be achieved synthetically, where an optimization process is employed to create the optimal input that maximizes the neuron’s activation [18, 19, 20], or naturally, by finding such inputs within a data corpus [21]. Activation Maximization has been employed for explaining latent representations of models [22, 23], including probabilistic models [24], detection of backdoor attacks [25] and spurious correlations [26]. However, one of the key limitations of this methodology lies in its inability to scale, as it relies on users to manually audit maximization signals.
@@
A more scalable alternative approach involves linking neurons with human-understandable concepts through textual descriptions. Network Dissection [11] (NetDissect) is a pioneering method in this field, associating convolutional neurons with a concept based on the Intersection over Union (IoU) of neuron activation maps and ground truth segmentation masks. Building on this, Compositional Explanations of Neurons (CompExp) [12] enhanced explanation detail by enabling the use of compositional concepts—i.e., concepts constructed with logical operators. MILAN [13] further expanded this by allowing for open-vocabulary explanations, permitting the generation of descriptions beyond predefined labels. INVERT [16] adopted a compositional concept approach, enabling explanations for general neuron types without the need for segmentation masks. It assigns compositional labels based on a neuron’s ability to distinguish concepts, using the Area Under the Receiver Operating Characteristic Curve (AUC). FALCON [14] and CLIP-Dissect [15] compute image-text similarity with a CLIP model [27] for the most activating images and their corresponding captions or concept sets. Each method defines its optimization criteria, lacking a unified consensus on what constitutes a good explanation. For detailed descriptions of the methods and their optimization objectives, please refer to Appendix A.1. An overview of the different techniques is illustrated in Table 1.
@@
While significant effort has been made towards developing approaches and tools for evaluating local explanations [28, 29, 30], there has been relatively limited focus on evaluating global methods, in particular neuron description methods. Currently, to the best of our knowledge, there is no unified approach that allows for benchmarking across models and explanation methods. In their respective papers, the INVERT and CLIP-Dissect explanation methods evaluated the accuracy of their explanations by comparing the generated neuron labels with ground truth descriptions provided for neurons in the output layer of a network. However, this evaluation is limited to output neurons and fixed labels only. CLIP-Dissect additionally evaluates the quality of explanations by computing the Cosine Similarity in a sentence embedding space between the ground truth class name for each neuron and the explanation generated by the method. FALCON employs a human study conducted on Amazon Mechanical Turk to evaluate the concepts generated by the method. Participants are tasked with selecting the best explanation for each target feature from a selection of explanation methods, considering a given set of highly and lowly activating images. MILAN evaluates the performance of neuron labeling methods relative to human annotations using BERTScores [31]. While human studies are generally beneficial, the conventional setup can be misleading and may fail to fully capture the intended evaluation criteria, introducing potential biases. Typically, annotators describe the images that most strongly activate a neuron, and these descriptions are then compared to an automatic explanation. However, this approach primarily evaluates the alignment with the most activating images rather than the accuracy of the explanation in describing the neuron’s function. Moreover, these highly activating images may not accurately represent the neuron’s overall behavior, as they only reflect the maximum tail of the distribution.
3 Method
@@
Table 1: Comparison of characteristics of neuron description methods. The columns (from left to right) represent the explanation method used, its textual output type (fixed-label, compositional, or open-vocabulary), the type of neuron targeted for analysis (convolutional, scalar, or predetermined), the target metric the method optimizes (IoU, WPMI, AUC, etc.), whether the method relies on auxiliary black-box models for finding or generating explanations (img2txt model, CLIP), and whether the explanation method is architecture-agnostic, meaning it can be applied to any CV model. For a more detailed description of each method, refer to Appendix A.1.
@@
NetDissect [11] fixed-label conv. IoU
CompExp [12] compositional conv. IoU
MILAN [13] open-vocabulary conv. WPMI img2txt model
FALCON [14] open-vocabulary predetermined avg. CLIP score CLIP
CLIP-Dissect [15] open-vocabulary scalar SoftWPMI CLIP
INVERT [16] compositional scalar AUC
@@ 3.1 Preliminaries
Consider a Deep Neural Network (DNN) represented by the function $g:\mathcal{X}\rightarrow\mathcal{Z},$ where $\mathcal{X}\subset\mathbb{R}^{h\times w\times c}$ denotes the input image domain and $\mathcal{Z}\subset\mathbb{R}^{l}$ represents the model’s output domain. We can view the model as a composition of two functions, $F:\mathcal{X}\rightarrow\mathcal{Y},$ and $L:\mathcal{Y}\rightarrow\mathcal{Z},$ such that $g=L\circ F.$ Here $\mathcal{Y}\subset\mathbb{R}^{d\times w^{*}\times h^{*}}$, where $d\in\mathbb{N}$ is the number of neurons in the layer, and $w^{*},h^{*}\in\mathbb{N}$ represent the width and height of the feature map, respectively. The function $F,$ which we refer to as the feature extractor, can be chosen based on the layer of the model we aim to inspect. This could be an existing layer within the model or a concept bottleneck layer [32]. We refer to the $i$-th neuron within the layer as $f_{i}(\bm{x})=F_{i}(\bm{x}):\mathcal{X}\rightarrow\mathbb{R}^{w^{*}\times h^{*}}.$ Within the scope of this paper, we refer to explanation method as an operator $\mathcal{E}$ that maps a neuron to the textual description $s=\mathcal{E}(f_{i})\in\mathcal{S},$ where $\mathcal{S}$ is a set of potential textual explanations. The specific set of explanations depends on the implementation of the particular method (see Appendix A.1).
@@ 3.2 CoSy: Evaluating Open-Vocabulary Explanations
We assume that a good textual explanation for a neuron should provide a human-understandable description of an input that strongly activates the neuron. However, modern methods for explaining the functional purpose of neurons often provide open-vocabulary textual explanations, complicating the quantitative collection of natural data that represents the explanation. To address this issue, CoSy utilizes recent advancements in generative models to synthesize data points that correspond to the textual explanation. The response of a neuron to a set of synthetic images is measured and compared to the neuron’s activation on a set of control natural images representing random concepts. This comparison allows for a quantitative evaluation of the alignment between the explanation and the target neuron.
The parameters of the proposed method include a control dataset $\mathbb{X}_{0}=\{\bm{x}^{0}_{1},\dots,\bm{x}^{0}_{n}\}\subset\mathcal{X},n\in\mathbb{N}$, which consists of natural images representing the concepts on which the model was originally trained. Additionally, it incorporates a generative model $p_{M}$ used for synthesizing images, along with a specified number of generated images $m\in\mathbb{N}$. The control dataset typically includes a balanced selection of validation class images. Given a neuron $f_{i}$ and explanation $s\in\mathcal{S}$, CoSy evaluates the alignment between the explanation and a neuron in three consecutive steps, which are illustrated in Figure 1.
@@
    $\begin{split}\mathbb{A}_{0}=\{\sigma(f_{i}(\bm{x}^{0}_{1})),\dots,\sigma(f_{i}(\bm{x}^{0}_{n}))\}\in\mathbb{R}^{n},\\ \mathbb{A}_{1}=\{\sigma(f_{i}(\bm{x}^{1}_{1})),\dots,\sigma(f_{i}(\bm{x}^{1}_{m}))\}\in\mathbb{R}^{m},\end{split}$ (1)
    where $\sigma:\mathbb{R}^{w^{*}\times h^{*}}\rightarrow\mathbb{R}$ is an aggregation function for multi-dimensional neurons. Within the scope of our paper, we use Average Pooling as aggregation function
    $\sigma(\bm{y})=\frac{1}{w^{*}h^{*}}\underset{k\in[1,w^{*}],l\in[1,h^{*}]}{{\sum\bm{y}_{k,l}},}\bm{y}\in\mathcal{Y}\subset\mathbb{R}^{w^{*}\times h^{*}}.$ (2)
@@
    $\Psi_{\text{MAD}}(\mathbb{A}_{0},\mathbb{A}_{1})=\frac{\frac{1}{m}\sum_{b\in\mathbb{A}_{1}}b-\frac{1}{n}\sum_{a\in\mathbb{A}_{0}}a}{\sqrt{\frac{1}{n-1}\sum_{a\in\mathbb{A}_{0}}(a-\bar{a})^{2}}},$ (4)
    with mean control activation $\bar{a}=\frac{1}{n}\sum_{a\in\mathbb{A}_{0}}a$.
@@
4 Sanity Checks
To ensure the reliability of our proposed evaluation measure, all steps within our framework need to be subject to sanity checks [33]. In this section, we analyze the following: (1) which generative models and prompts provide the best similarity to natural images, (2) whether the model’s behavior on synthetic and natural images differs for the same class, and (3) validating that CoSy provides appropriate evaluation scores for true and random explanations, given a known ground truth class for the neuron.
@@ 4.1 Synthetic Image Reliability
One of the key features of CoSy is its reliance on generative models to translate textual explanations of neurons into the visual domain. Thus, it is essential that the generated images reliably resemble the textual concepts. In the following section, we present an experiment where we varied several parameters of the generation procedure and evaluated the visual similarity between generated images and natural ones, focusing on concepts for which we have a collection of natural images.
For our analysis, we used only open-source and freely available text-to-image models, namely Stable Diffusion XL 1.0-base (SDXL) [34] and Stable Cascade (SC) [35]. We also varied the prompts for image generation. To measure the similarity between synthetic images and natural images corresponding to the same concept, we used Cosine Similarity (CS) in the CLIP embedding space with the CLIP-ViT-B/32 model [27]. We select a set of 50 random concepts from the 1,000 classes in the ImageNet validation dataset [36]. For each [concept] we use five different prompts and employ them with SDXL and SC models, generating 50 images per class. We then measure the CS between image pairs of the same class.
Figure 2: An overview of the impact of varying the prompt on the similarity between natural and synthetic images, using two text-to-image models. Left: average Cosine Similarity (CS) across all natural and synthetic images over all classes are reported. Higher CS values are better, indicating greater similarity between the images. Right: an illustration of the visual differences produced by the SDXL and SC models in response to diverse prompts for the concept “submarine”, and natural images from the ImageNet validation dataset [36]. Our results show that both SDXL and SC generate similar images, with SDXL generally being more closely aligned with natural images than SC.
Figure 2 illustrates the comparison across all generative models and prompts in terms of CS of generated images to natural images of the same class. The results indicate that when using Prompt 5 as input to SDXL, the synthetic images show the highest similarity to natural images. The performance is generally best with the most detailed prompt (5) and closely aligns with prompts 1, 3, and 4. Moreover, SDXL appears to be slightly more effectively realizing detailed prompts than SC. As anticipated, the poorly constructed prompt (2) results in the lowest similarity to natural images for both models. To address prompt bias and dataset dependency, we compare the object-focused ImageNet with the scene-focused Places365 (see Appendix A.3). We find that close-up prompts work well for object-centric datasets, while general prompts like “photo of” are better for scene-based datasets. If not stated otherwise, for all following experiments, Prompt 5 together with SDXL model was employed for image generation.